Calculate the maximum strength for the following two cases, using the Griffith equation. Given: (i) mineral platelets have thickness of 1 mm and diameter of 10 nm;

(ii) mineral platelets have thickness of 1 nm and diameter of 50 nm.

Assume γsurf = 1 J/m2 ; EHAP = 100 GPa.


1
Expert's answer
2021-08-16T02:20:50-0400

Solution;

Fracture strength of mineral platelets by Griffith criteria is given by;

"\sigma_m^f=\alpha E_m\Psi"

In which;

"\Psi=\sqrt{\frac{\gamma}{E_mh}}"

"\alpha\approx\sqrtπ"

"\gamma" is the surface energy;

h is diameter of the mineral platelet.

(i)

"\sigma_m^f" ="\sqrtπ"×100×109×"\sqrt{\frac{1}{100×10^9×10×10^{-9}}}"

"\sigma_m^f=5.605GPa"

(ii)

"\sigma_m^f=\sqrtπ" ×100×109×"\sqrt{\frac{1}{50×100}}"

"\sigma_m^f=2.507GPa"